# Sharp bounds on the half-space two-point function for high-dimensional Bernoulli percolation

### v1

@@ 1. $(a)$
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\leq C\frac{1+m}{1+|x-y|^{d-1}}.$ (1.6)
@@ 2. $(b)$
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\geq c\frac{1+m}{1+|x-y|^{d-1}}.$ (1.7)
@@
Finally, let us mention that the upper bound in Theorem 1.4 is also useful in the recent [ASS25].
@@
where we used that $\min(x_{1},n)\lesssim r_{x,y}$ and $\min(y_{1},n)\lesssim r_{y,x}$. In the situation where $n=0$ (resp. $n<L$ in the spread-out case), a similar bound holds trivially (to the cost of potentially increasing $C_{1}$).
@@
which strengthens the result of Proposition 2.1.
@@
Using (1.6) and (2.5) for $x_{1}=0$ (resp. $x_{1}\leq L-1$), we deduce that, for every $x\in\mathbb{H}$ and $u\in\partial B_{n}(x)$,
@@
It remains to treat the case $n<\tfrac{1}{2}x_{1}$. Observe that for this choice, one has $B_{n}(x)=\{u\in\mathbb{Z}^{d}:|u-x|\leq n\}$ (that is, the $\mathbb{Z}^{d}$-box of radius $n$ around $x$ is fully included in $\mathbb{H}$). Thus, we may use [DCT16] to conclude that, for any $\varepsilon\in(0,1/2)$,
@@
The proof of Proposition 2.1 is technically more involved. We will need a number of preliminary results and notations. In Section 4.1, we present a general lower bound for the probability that two sets are connected, which we express in terms of their (d4)(d-4)-capacity. In Section 4.2, we recall all the necessary definitions to introduce the theory of regularity of [KN11], taking here the viewpoint of [ASS25]. Finally, we give the proof of Proposition 2.1 in Section 4.3. Recall that we have assumed that $d>6$ and that (*1) holds.
@@ 4.2 Regular points, line good points, extended cluster
$\mathcal{T}_{s}(z):=\{|\mathcal{C}(z;B_{n}(x))\cap B(z,s)|\leq s^{4}(\log s)^{7}\}\\ \cap\{|\mathcal{C}(z;B_{n}(x))\cap B(z,s)\cap\partial B_{n}(x)|\leq s^{2}(\log s)^{7}\}.$ (4.19)
@@
We defer the proof of this lemma to Section 4.4 and introduce now the notion of $K$-line good points. For this, one first needs to consider a maximal subset of the set $X_{n}^{\varepsilon,K-\textup{reg}}(x)$ of $K$-regular points, which has the property that all its points are at distance at least $2K$ one from each other. Denote by $\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$ one such maximal subset chosen uniformly at random. Then, if $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we consider a line segment of length $K$ emanating from $z$, outside $B_{n}(x)$ and orthogonal to its boundary (choose one arbitrarily if there are many). Call $z^{\prime}$ the endpoint of this line segment. We say that $z^{\prime}$ is a $K$-line good point if all the edges on the line segment between $z$ and $z^{\prime}$ are open. More generally, for any $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we denote by $L_{z}$ the maximal open segment emanating from $z$ orthogonally to $B_{n}(x)$, of length at most $K$.
@@
Now, by definition of a $K$-regular point, one has for any fixed $a\in\partial_{*}A$,
$\sum_{a^{\prime}\in\partial_{*}A}(1+\|a-a^{\prime}\|)^{4-d}\lesssim\sum_{i\geq\log_{2}(K)}\frac{2^{2i}\cdot i^{7}}{2^{(d-4)i}}\lesssim 1,$ (4.23)
where the implicit constants do not depend on $x$, $n$, $K$, $\varepsilon$, and $A$. This concludes the proof. ∎
@@ 4.3 Conclusion
$\begin{split}\mathbb{P}\big[a\xleftrightarrow{\textrm{via }H\>}b\big]&\leq\sum_{u\in H}\tau_{\mathbb{H}}(a,u)\tau_{\mathbb{H}}(u,b)\lesssim\frac{1}{n^{d-2}}\Big(\sum_{u\in H\cap A}\tau_{\mathbb{H}}(a,u)+\sum_{u\in H\cap B}\tau_{\mathbb{H}}(u,b)\Big)\\ &\lesssim\frac{1}{n^{d-2}}\sum_{i\geq\log_{2}(K)}\frac{2^{4i}\cdot i^{7}}{2^{(d-2)i}}\lesssim\frac{1}{n^{d-2}\sqrt{K}},\end{split}$ (4.32)
which gives (4.31)—and hence (4.30)—by choosing $K$ large enough.
@@
The interest of this event, when compared to $\mathcal{T}_{s}(z)$, is that it only depends on the configuration of the percolation inside the box $B(z,s^{d})$, and is thus a purely local event, while to determine whether $\mathcal{T}_{s}(z)$ holds or not, one needs a priori to know the configuration in the whole box $B_{n}(x)$. The drawback is that this event is a priori less likely than $\mathcal{T}_{s}(z)$, but as Lemma 4.6 below shows, it is still extremely likely, and furthermore, the following simple fact holds by construction (see Claim 4.1 in [KN11] or Claim 5.4 in [ASS25]).
@@
We now fix $s\geq K$ such that $8s^{d}\leq n$ and consider the set $U=\{u\in\mathbb{Z}^{d}:u_{i}\in\{0,s^{d}\}\ \forall i=1,\dots,d\}$. For each $w\in U$, we define
$\mathcal{B}(w)=\{B(z,2s^{d}):z\in w+4s^{d}\cdot\mathbb{Z}^{d}\}.$ (4.45)
Denote by $Q(w)$ the union of all the boxes of this partition which intersect $\partial B_{n}^{\varepsilon}(x)$, and let $N(w)$ be the number of those which are connected to $x$ inside $B_{n}(x)\setminus Q(w)$. By Lemma 4.7 and the remark following it, one has $c_{2},C_{2}>0$ such that, for any $w\in U$,

### v2

@@ 1. $(a)$
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\leq C\frac{1+m}{1+|x-m\mathbf{e}_{1}|^{d-1}}.$ (1.6)
@@ 2. $(b)$
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\geq c\frac{1+m}{1+|x-m\mathbf{e}_{1}|^{d-1}}.$ (1.7)
@@
Finally, let us mention that the upper bound in Theorem 1.4 is also useful in the recent [ASS25] (see Lemma 4.3 there).
@@
where (recall that $r_{x,y}=\min(x_{1},|x-y|)$) we used that $\min(x_{1},n)\lesssim r_{x,y}$ and $\min(y_{1},n)\lesssim r_{y,x}$. In the situation where $n=0$ (resp. $n<L$ in the spread-out case), a similar bound holds trivially (to the cost of potentially increasing $C_{1}$), since $\max(r_{x,y},r_{y,x})\leq\max(1,L)$.
@@
which strengthens the result of Proposition 2.1. Let us provide a short proof of (2.14). For $u,v$ as above, (*1) gives that $\tau_{\mathbb{H}}(u,v)\lesssim\frac{1}{|x-y|^{d-2}}=\frac{1}{n^{d-2}}$, and (2.5) gives that
$\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\lesssim\frac{1+\min(x_{1},n)}{n},\qquad\sum_{v\in\partial B_{n}(y)}\tau_{B_{n}(y)}(y,v)\lesssim\frac{1+\min(y_{1},n)}{n}.$ (2.15)
Combining these observations, we get that the sum on the right-hand side of (2.14) is bounded by a quantity of order $\frac{(1+\min(x_{1},n))(1+\min(y_{1},n))}{n^{d}}$, which is itself bounded by (a multiple of) $\tau_{\mathbb{H}}(x,y)$ by Theorem 1.4.
@@
Write $x=(x_{1},x_{\bot})$, with $x_{\bot}\in\mathbb{Z}^{d-1}$. For every $v=(v_{1},v_{\bot})\in\partial B_{n/4}(u)\cap B_{n}(x)$, letting $\tilde{x}=x-(0,v_{\bot})$ and $\tilde{v}=v-(0,x_{\bot})$, (1.6) gives that
$\mathbb{P}[x\xleftrightarrow{B_{n}(x)\>}v]=\mathbb{P}[\tilde{x}\xleftrightarrow{B_{n}(\tilde{x})\>}v_{1}\mathbf{e}_{1}]=\mathbb{P}[\tilde{v}\xleftrightarrow{B_{n}(x_{1}\mathbf{e}_{1})\>}x_{1}\mathbf{e}_{1}]\lesssim\frac{\min(1+v_{1},1+x_{1})}{1+|x-v|^{d-1}}\lesssim\frac{1+\min(x_{1},n)}{n^{d-1}},$ (3.2)
where we used translation invariance in the first two equalities. Plugging (3.2) in (3.1) and using (2.5) for $x_{1}=0$ (resp. $x_{1}\leq L-1$), we deduce that, for every $x\in\mathbb{H}$ and $u\in\partial B_{n}(x)$,
@@
It remains to treat the case $n<\tfrac{1}{2}x_{1}$. Observe that for this choice, one has $B_{n}(x)=\{u\in\mathbb{Z}^{d}:|u-x|\leq n\}$ (that is, the $\mathbb{Z}^{d}$-box of radius $n$ around $x$ is fully included in $\mathbb{H}$). Thus, we may use [DCT16] (which gives that $\varphi_{p_{c}}(B_{n}(x)):=\sum_{\begin{subarray}{c}u\in\partial B_{n}(x)\\ v\notin B_{n}(x)\\ u\sim v\end{subarray}}\tau_{B_{n}(x)}(x,u)p_{c}\geq 1$) to conclude that, for any $\varepsilon\in(0,1/2)$,
@@
The proof of Proposition 2.1 is technically more involved. We will need a number of preliminary results and notations.
We rely on the notion of regular points initially introduced in [KN11], and recently revisited in [ASS25]. In particular, we take advantage of the geometric and convenient definition of regular points chosen in [ASS25]. Most of the arguments below are adaptations of intermediate results that already appeared in this paper. It is worth mentioning that our setting is often simpler. For instance, our Lemma 4.3 requires an averaged estimate on the expected number of regular points, while [ASS25, Section 6] derives finer and stronger estimates (quite similarly to what is done in [KN11]). Our main contribution is to use these regular points to derive “reversed Simon–Lieb type inequalities”.
In Section 4.1, we present a general lower bound for the probability that two sets are connected, which we express in terms of their (d4)(d-4)-capacity. In Section 4.2, we recall all the necessary definitions to introduce the theory of regularity of [KN11], taking here the viewpoint of [ASS25]. Finally, we give the proof of Proposition 2.1 in Section 4.3. Recall that we have assumed that $d>6$ and that (*1) holds.
@@ 4.2 Regular points, line good points, extended cluster
$\mathcal{T}_{s}(z):=\{|\mathcal{C}(z;B_{n}(x))\cap B_{s}(z)|\leq s^{4}(\log s)^{7}\}\\ \cap\{|\mathcal{C}(z;B_{n}(x))\cap B_{s}(z)\cap\partial B_{n}(x)|\leq s^{2}(\log s)^{7}\}.$ (4.19)
@@
We defer the proof of this lemma to Section 4.4 and introduce now the notion of $K$-line good points. For this, one first needs to consider a maximal subset of the set of $K$-regular points of $\partial B_{n}^{\varepsilon}(x)$ which has the property that all its points are at distance at least $2K$ one from each other. Denote by $\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$ one such maximal subset chosen uniformly at random. Then, if $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we consider a line segment of length $K$ emanating from $z$, outside $B_{n}(x)$ and orthogonal to its boundary (choose one arbitrarily if there are many). Call $z^{\prime}$ the endpoint of this line segment. We say that $z^{\prime}$ is a $K$-line good point if all the edges on the line segment between $z$ and $z^{\prime}$ are open. More generally, for any $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we denote by $L_{z}$ the maximal open segment emanating from $z$ orthogonally to $B_{n}(x)$, of length at most $K$.
@@
Let $\omega$ be a percolation configuration realizing $\{\mathcal{C}_{n}^{e}(x)=A\}$ (note that it exists since $A$ is $K$-admissible for the pair $(x,n)$). By definition, to every fixed $a\in\partial_{*}A$ corresponds a unique $z_{a}\in\partial B_{n}(x)$ such that $\omega$ realises the events $\{z_{a}\xleftrightarrow{B_{n}(x)\>}x\}$ and $\mathcal{T}_{s}(z_{a})$ for every $s\geq K$. Thus, one has that, for every $a\in\partial_{*}A$, and every $s\geq K$,
$|\partial_{*}A\cap B_{s}(a)|\lesssim|\partial_{*}A\cap B_{s}(z_{a})|\lesssim|\mathcal{C}(z_{a};B_{n}(x))\cap B_{s}(z_{a})\cap\partial B_{n}(x)|\leq s^{2}(\log s)^{7},$ (4.23)
where the implicit constants do not depend on $x$, $n$, $K$, $\varepsilon$, and $A$. As a consequence, for every fixed $a\in\partial_{*}A$, we find that
$\sum_{a^{\prime}\in\partial_{*}A}(1+\|a-a^{\prime}\|)^{4-d}\lesssim 1+\sum_{i\geq\log_{2}(K)}\frac{|\partial_{*}A\cap(B_{2^{i+1}}(a)\setminus B_{2^{i}}(a))|}{2^{i(d-4)}}\lesssim 1+\sum_{i\geq\log_{2}(K)}\frac{2^{2i}(\log 2^{i})^{7}}{2^{i(d-4)}}\lesssim 1,$ (4.24)
where in the first inequality we used that (by definition) $(\partial_{*}A\setminus\{a\})\cap B_{K}(a)=\emptyset$, in the second inequality we used (4.23), and where (again) the implicit constants do not depend on $x$, $n$, $K$, $\varepsilon$, and $A$. This concludes the proof. ∎
@@ 4.3 Conclusion
$\begin{split}\mathbb{P}\big[a\xleftrightarrow{\textrm{via }H\>}b\big]&\leq\sum_{u\in H}\tau_{\mathbb{H}}(a,u)\tau_{\mathbb{H}}(u,b)\lesssim\frac{1}{n^{d-2}}\Big(\sum_{u\in H\cap A}\tau_{\mathbb{H}}(a,u)+\sum_{u\in H\cap B}\tau_{\mathbb{H}}(u,b)\Big)\\ &\lesssim\frac{1}{n^{d-2}}\sum_{i\geq\log_{2}(K)}\frac{2^{4i}(\log 2^{i})^{7}}{2^{i(d-2)}}\lesssim\frac{1}{n^{d-2}\sqrt{K}},\end{split}$ (4.33)
where we used that
$\sum_{u\in H\cap A}\tau_{\mathbb{H}}(a,u)\lesssim\sum_{i\geq\log_{2}(K)}\frac{|H\cap A\cap(B_{2^{i+1}}(a)\setminus B_{2^{i}-1}(a))|}{2^{i(d-2)}}\lesssim\sum_{i\geq\log_{2}(K)}\frac{2^{4i}(\log 2^{i})^{7}}{2^{i(d-2)}},$ (4.34)
which follows by using that $H\cap B_{K-1}(a)=\emptyset$, and by proving—similarly to (4.23)—that $|H\cap A\cap B_{s}(a)|\lesssim s^{4}(\log s)^{7}$ for every $s\geq K$. As a consequence, (4.33) gives (4.32)—and hence (4.31)—by choosing $K$ large enough.
@@
The interest of this event, when compared to $\mathcal{T}_{s}(z)$, is that it only depends on the configuration of the percolation inside the box $B_{s^{d}}(z)$, and is thus a purely local event, while to determine whether $\mathcal{T}_{s}(z)$ holds or not, one needs a priori to know the configuration in the whole box $B_{n}(x)$. The drawback is that this event is a priori less likely than $\mathcal{T}_{s}(z)$, but as Lemma 4.6 below shows, it is still extremely likely, and furthermore, the following simple fact holds by construction (see Claim 4.1 in [KN11] or Claim 5.4 in [ASS25]).
@@
We now fix $s\geq K$ such that $8s^{d}\leq n$ and consider the set $U=\{u\in\mathbb{Z}^{d}:u_{i}\in\{0,s^{d}\},\ \forall i=1,\dots,d\}$. For each $w\in U$, we define
$\mathcal{B}(w)=\{B_{2s^{d}}(z):z\in w+4s^{d}\cdot\mathbb{Z}^{d}\}.$ (4.47)
Denote by $Q(w)$ the set of all the boxes of this partition which intersect $\partial B_{n}^{\varepsilon}(x)$. We explore them using the following algorithm. First, we reveal the whole percolation configuration outside the union of these boxes. Next, if at least one of the boxes of $Q(w)$ is connected to $x$ by an open path, that remains in the explored region, we choose one at random and reveal the configuration inside it, and we continue as long as there still exist an unexplored box of $Q(w)$ which is connected to $x$ via an open path in the explored region. We let $N(w)$ be the number of boxes of $Q(w)$ which have been revealed by this algorithm. Note that during this exploration procedure, each time we reveal the configuration inside a new box, almost surely, conditionally on the configuration outside the box, the probability that it contains a pioneer point is at least $c_{2}\exp(-C_{2}(\log s)^{2})$ by Lemma 4.7 and the remark following it, for some constant $c_{2},C_{2}>0$. Hence, for any $w\in U$,